OpenStudy (anonymous):

37 1/2% of what number is 27

OpenStudy (sbuck98):

27* 37.5%=? 10.125 which is rounded to 10.13 So 27-10.13=16.87 Do you understand it? If not I'll help more.

OpenStudy (anonymous):

yes bc that is not one of my answers

OpenStudy (jdoe0001):

well, THAT number is ... the 100%, whatever that is and 27 is 37 1/2% of that so \(\begin{array}{ccllll} value&\% \\\hline\\ 27&37\frac{1}{2}\\ x&100 \end{array}\implies \cfrac{27}{x}=\cfrac{37\frac{1}{2}}{100}\) solve for "x"